OpenStudy (anonymous):

the line parallel to x-3=-12 that passes through (-3,4) Any Help?

OpenStudy (anonymous):

it must have same slope

OpenStudy (anonymous):

what is the slope though

OpenStudy (anonymous):

how would you find the slope?

OpenStudy (anonymous):

substitute (-3, 4)>> (x1,y1) in:Eq of the line: (y-y1)=m(x-x1),m:slope

OpenStudy (anonymous):

???

OpenStudy (anonymous):

is the question complete?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

are you sure ?

OpenStudy (anonymous):

yes, all it says is: The line parallel to x-3y=-12 that passes through (-3,4)

OpenStudy (anonymous):

...

OpenStudy (anonymous):

y-4=(1/3)(x-0) the slope is (1/3)

OpenStudy (anonymous):

ok thanks!

OpenStudy (anonymous):

substitute (-3, 4)>> (x1,y1) in:: (y-y1)=(1/3)(x-x1)